# On the Expressive Power of a Variant of the Looped Transformer

### v1

@@ Abstract
Besides natural language processing, transformers exhibit extraordinary performance in solving broader applications, including scientific computing and computer vision. Previous works try to explain this from the expressive power and capability perspectives that standard transformers are capable of performing some algorithms. To empower transformers with algorithmic capabilities and motivated by the recently proposed looped transformer (Yang et al., 2024; Giannou et al., 2023), we design a novel transformer block, dubbed Algorithm Transformer (abbreviated as AlgoFormer). Compared with the standard transformer and vanilla looped transformer, the proposed AlgoFormer can achieve significantly higher expressiveness in algorithm representation when using the same number of parameters. In particular, inspired by the structure of human-designed learning algorithms, our transformer block consists of a pre-transformer that is responsible for task pre-processing, a looped transformer for iterative optimization algorithms, and a post-transformer for producing the desired results after post-processing. We provide theoretical evidence of the expressive power of the AlgoFormer in solving some challenging problems, mirroring human-designed algorithms. Furthermore, some theoretical and empirical results are presented to show that the designed transformer has the potential to be smarter than human-designed algorithms. Experimental results demonstrate the empirical superiority of the proposed transformer in that it outperforms the standard transformer and vanilla looped transformer in some challenging tasks.
Machine Learning, ICML
@@ 1 Introduction
The emergence of the transformer architecture (Vaswani et al., 2017) marks the onset of a new era in natural language processing. Transformer-based large language models (LLMs), such as BERT (Devlin et al., 2019) and GPT-3 (Brown et al., 2020), revolutionized impactful language-centric applications, including language translation (Vaswani et al., 2017; Raffel et al., 2020), text completion/generation (Radford et al., 2019; Brown et al., 2020), sentiment analysis (Devlin et al., 2019), and mathematical reasoning (Imani et al., 2023; Yu et al., 2024; Zheng et al., 2023). Beyond the initial surge in LLMs, these transformer-based models have found extensive applications in diverse domains such as computer vision (Dosovitskiy et al., 2021), time series (Li et al., 2019), bioinformatics (Zhang et al., 2023b), and addressing various physical problems (Cao, 2021). While many studies have concentrated on employing transformer-based models to tackle challenging real-world tasks, yielding superior performances compared to earlier models, the mathematical understanding of transformers remains incomplete.
The initial line of research interprets transformers as function approximators. Edelman et al. (2022) delve into the learnability and complexity of transformers for the class of sparse Boolean functions. Gurevych et al. (2022) explore the binary classification tasks, demonstrating that the transformer can circumvent the curse of dimensionality with a suitable hierarchical composition model for the posterior probability. Takakura & Suzuki (2023) investigate the approximation ability of transformers as sequence-to-sequence functions with infinite-dimensional inputs and outputs. They highlight that the transformer can avoid the curse of dimensionality under the smoothness conditions, due to the feature extraction and property sharing mechanism. However, viewing transformers solely as functional approximators, akin to feed-forward neural networks and convolutional neural networks, may not provide a fully satisfying understanding.
@@
Another line of studies focuses on understanding transformers as algorithm learners. Garg et al. (2022) empirically investigate the performance of transformers in in-context learning, where the input tokens are input-label pairs generated from classical machine learning models, e.g., (sparse) linear regression and decision tree. They find that transformers can perform comparably as standard human-designed machine learning algorithms. Some subsequent works try to explain the phenomenon. Akyürek et al. (2023) characterize decoder-based transformer as employing stochastic gradient descent for linear regression. Bai et al. (2023) demonstrate that transformers can address statistical learning problems and employ algorithm selection, such as ridge regression, Lasso, and classification problems. Zhang et al. (2023a) and Huang et al. (2023) simplify the transformer model with reduced active parameters, yet reveal that the simplified transformer retains sufficient expressiveness for in-context linear regression problems. In Ahn et al. (2023), transformers are extended to implement preconditioned gradient descent. The looped transformer is proposed in (Giannou et al., 2023), and is shown to have the potential to perform basic operations (e.g., addition and multiplication), as well as implicitly learn iterative algorithms  (Yang et al., 2024). More related and interesting studies can be found in (Huang et al., 2023; Von Oswald et al., 2023; Mahankali et al., 2024; Sun et al., 2023).
@@
In this paper, inspired by the recently proposed looped transformer (Yang et al., 2024; Giannou et al., 2023), we propose a novel transformer block, which we refer to as AlgoFormer, and strictly enforce it as an algorithm learner by regularizing its architecture. The transformer block consists of three sub-transformers, i.e., the pre-, looped, and post-transformers, designed to perform distinct roles. The pre-transformer is responsible for preprocessing the input data, and formulating it into some mathematical problems. The looped transformer acts as an iterative algorithm in solving the hidden problems. Finally, the post-transformer handles suitable postprocessing to produce the desired results. In contrast to standard transformers, the AlgoFormer is more likely to implement algorithms, due to its algorithmic structures shown in Figure 1.
Our main contributions can be summarized as follows:
  • We introduce a novel transformer block (as shown in Figure 1), inspired by looped transformer (Yang et al., 2024; Giannou et al., 2023), namely the AlgoFormer. This block is designed as algorithm learners, mimicking the structure of human-designed algorithms. Compared with the standard transformer, AlgoFormer enjoys a lower parameter size but higher expressiveness in specific algorithmic learning tasks; Compared with the vanilla looped transformer, AlgoFormer can represent complex algorithms more efficiently; see Section 2.
  • We theoretically show the expressive power of the AlgoFormer in solving three challenging tasks in implementing some practical human-designed algorithms (Theorems 3.1-3.3), including some task-specific pre- and post-processing, as well as the gradient descent; see Section 3.
  • Beyond the gradient descent, we prove that the AlgoFormer can implement the (second-order) Newton’s method in linear regression problems (Theorem 4.1). While our results primarily focus on encoder-based transformers, we also extend our findings to decoder-based transformers (Theorem 4.2); see Section 4.
  • We experimentally investigate the behavior of the AlgoFormer with different hyperparameters in handling some challenging in-context learning tasks. The empirical results support the stronger expressiveness and powerfulness of the AlgoFormer, compared with the standard transformer and vanilla looped transformer; see Section 5.
This paper is organized as follows. The motivation for the design of the transformer block, including a brief introduction to transformer layer architectures, the algorithmic structure of the proposed transformer block, and the advantages of lower parameter size and higher expressiveness, are presented in Section 2. We provide detailed results for the expressiveness of the designed AlgoFormer in tackling three challenging tasks, in Section 3. In Section 4, we additionally show that the designed transformer is capable of implementing Newton’s method, a second-order optimization algorithm, beyond the gradient descent. We also extend our results for decoder-only transformers, which can only access data in the previous tokens for regression. In Section 5, experimental results studying the behavior of the designed transformer block are reported. We also empirically compare it with the standard transformer and vanilla looped transformer. Some concluding remarks and potential works are discussed in Section 6.
2 Motivation
@@
In this section, we mainly discuss the construction and intuition of the AlgoFormer. Its advantages over standard transformer is then conveyed. Before going into details, we first elaborate the mathematical definition of transformer layers.
@@ 2.1 Preliminaries
A One-layer transformer is mathematically formulated as
@@ 2.2 Algorithmic Structures of Transformers
Figure 1: Algorithmic structure of the AlgoFormer. The pre-transformer conducts “statements 1” in the Algorithm; the looped transformer performs “statements 2” inner the for loop; the post-transformer carries out “statements 3”. Here, $\text{TF}_{\text{pre}}$, $\text{TF}_{\text{loop}}$, and $\text{TF}_{\text{post}}$ are multi-layer transformers; “statements” represent some fundamental operations in classical algorithms.
As discussed in the introduction, rather than simply interpreting it as a function approximator, the transformer may in-context execute some implicit algorithms learned from training data. However, it is still unverified that the standard multi-layer transformer is exactly performing algorithms.
Looped transformer. Although those works theoretically explain the success of transformers in solving (sparse) linear regression problems by the realization of (proximal) gradient descent, very little evidence implies that standard transformers are exactly implementing optimization algorithms. Note that the same mode of computation is conducted in each step of iteration. Equipped with the supposition that transformers can perform some basic operations, a shallow transformer is sufficient to represent iterative algorithms, as shown in the green part of Figure 1. The idea is first mentioned in Giannou et al. (2023), where they show that transformers can implement matrix addition and multiplication. Therefore, some iterative algorithms in scientific computing can be potentially realized by looped transformers. However, their construction does not explicitly imply the expressive power of transformers defined in Equation (1). Recently, Yang et al. (2024) empirically studied the behavior of looped transformers with different choices of hyperparameters in solving regression problems. Their experimental results further confirm the hypothesis that transformers are exactly learning iterative algorithms in solving linear regression problems, by strictly regularizing the transformer structure as looped transformers.
AlgoFormer. As shown in the green part of Figure 1, vanilla looped transformers (Yang et al., 2024; Giannou et al., 2023) admit the same structure as iterative algorithms, whose expressive power may be limited. However, real applications are usually much more complicated than linear regression problems. For example, given a task with data pairs, a well-trained researcher may first pre-process the data under some prior knowledge, and then formulate a mathematical (optimization) problem. Following that, some designed solvers, usually iterative algorithms, are performed. Finally, the desired results are obtained after further computation, namely, post-processing. The designed AlgoFormer (Algorithm Transformer), visualized in Figure 1, enjoys the same structure as wide classes of algorithms. Specifically, we separate the transformer block into three parts, i.e., pre-transformer $\text{TF}_{\text{pre}}$, looped transformer $\text{TF}_{\text{loop}}$, and post-transformer $\text{TF}_{\text{post}}$. Here, those three sub-transformers are standard multi-layer transformers in Equation (1). Given the input token vectors $\bm{X}$ and the number of iteration steps $T$, the output admits
@@
Compared with standard transformers, the AlgoFormer acts more as the algorithm learner, by strictly regularizing the loop structure. In comparison to the results presented in the previous work Giannou et al. (2023), our construction of AlgoFormer is closer to the transformer in real applications. In their approach, the design of the looped transformer necessitates task-specific knowledge, involving operations like token order switching. In contrast, our model is task-independent, emphasizing the learning of algorithms solely from data rather than relying on potentially unknown prior knowledge. In contrast to the looped transformer in Yang et al. (2024), we introduce pre- and post-transformers, which play essential roles in real applications, and the designed transformer block can represent complex algorithms and solve challenging tasks more efficiently.
Diffusion model. The looped transformer exhibits a structure similar to diffusion models (Sohl-Dickstein et al., 2015). This resemblance arises from the fact that diffusion models generate a clean image from a given noisy image, following a specific dynamical system governed by stochastic partial differential equations. It’s noteworthy that (iterative) optimization algorithms can also be viewed as a kind of dynamical system. Several related works leverage diffusion models for generating model parameters as meta-learners, simulating optimization dynamics (Krishnamoorthy et al., 2023; Peebles et al., 2022).
Meta-learner. The concept of a transformer as an algorithm learner shares connections with meta-learners like MAML (Finn et al., 2017) and Meta-SGD (Li et al., 2017). Meta-learners provide model parameters as initialization in a way that a small number of parameter updates and a limited amount of training samples are sufficient for achieving strong generalization performance on new tasks. Similarly, the AlgoFormer exhibits the capability to address new tasks with a modest number of (in-context) training samples.
2.3 Empirical Advantages
Figure 2: The validation error of trained models (the standard transformer and the AlgoFormer), assessed on sparse linear regression tasks. By choosing suitable hyperparameters (for example, we set $(T,\Delta T)=(12,10)$), defined in Equation (6), the proposed transformer has better performance than the standard transformer on the sparse linear regression.
In this subsection, we experimentally compare the performance of the AlgoFormer with the standard 12-layer transformer (i.e., GPT-2 model) in solving sparse linear regression. Note that from the algorithm learning perspective, the standard GPT-2 model can merely conduct at most 12 steps of computation, while our model exhibits higher expressive power with much lower parameter sizes but deeper loops. It is a significant advantage of the AlgoFormer, which enjoys algorithmic structures.
Lower parameter size. In the experiment, we adopt the transformer block where the three sub-transformers are all one-layer. Compared with the standard transformer (e.g., the GPT-2 model), our model is 4 times smaller in terms of the parameter size. It is imperative to note that our primary motivation was not centered around reducing the parameter size through weight sharing. Instead, our focus lies in enhancing the transformer’s role as an algorithm learner by incorporating algorithmic structures.
Higher expressiveness. Considering the (iterative) algorithm learning, in theory, the designed AlgoFormer admits even higher expressiveness than the GPT-2 model, although the latter enjoys more flexibility (large parameter size). It is also validated from the experiments (shown in Figure 2) that the designed transformer performs better than the standard transformer in sparse regression problems. The hard regularization on the transformer structure moves it closer to the algorithm learner. Note that the expressiveness of general weight-sharing models is often diminished due to their lower number of flexible parameters. However, the AlgoFormer demonstrates higher expressiveness, especially in specific iterative algorithm learning tasks, where the incorporation of deep loops plays a crucial role. In contrast to the vanilla looped transformer, the introduction of pre- and post-transformer components enables the transformer to learn more complicated and broader classes of algorithms.
Structure regularization. The transformer is intentionally designed with a structure similar to classical algorithms. The structure regularization of the transformer forces it to learn and conduct some implicit algorithms in solving in-context problems. In essence, it transforms the model into an explainable machine learning model, moving away from being a completely opaque, data-centric black-box model.
We conduct in-context learning of transformers on sparse linear regression problems. We compare the performances of the AlgoFormer (with different hyperparameters) and the GPT-2 model. As shown in Figure 2, for both two models, more in-context samples contribute to lower error, consistent with our intuition. In those simple examples, the designed AlgoFormer exhibits more competitive capabilities with the standard transformer. Due to the space limits, we put the details of the experiments in Appendix B.1. Further experiments on various tasks are presented in Section 5.
@@ 3 Expressive Power
In this section, we theoretically show by construction that AlgoFormer is capable of solving some challenging tasks, akin to human-designed algorithms. The core idea is as follows. Initially, the pre-transformer undertakes the crucial task of preprocessing the input data, such as representation transformation. The looped transformer is responsible for iterative algorithms in solving optimization problems. Finally, it is ready to output the desired result by the post-transformer. Through the analysis of AlgoFormer’s expressive power in addressing these tasks, we expect its potential to make contributions to the communities of scientific computing and machine learning.
@@ 3.1 Regression with Representation
We consider regression problems with representation, where the output behaves as a linear function of the input with a fixed representation function. Here, we adopt the $L$-layer MLPs with (leaky) ReLU activation function as the representation function $\Phi^{*}(\cdot)$. Specifically, we generate each in-context sample by first sampling the linear weight $\bm{A}$ from the prior $\mathcal{P}_{A}$, and then generating the input-label pair $\{(\bm{x}_{i},\bm{y}_{i})\}$ with $\bm{x}_{i}\in\mathbb{R}^{d}\sim\mathcal{P}_{x}$, $\bm{y}_{i}=\bm{A}\Phi^{*}(\bm{x}_{i})+\bm{\epsilon}_{i}$ and $\bm{\epsilon}_{i}\sim\mathcal{N}\left(\bm{0},\sigma^{2}\bm{I}\right)$. We aim to find the test label $\bm{y}_{\text{test}}:=\bm{A}\Phi^{*}(\bm{x}_{\text{test}})$, given the in-context samples and test data $\{\bm{x}_{1},\bm{y}_{1},\cdots,\bm{x}_{N},\bm{y}_{N},\bm{x}_{\text{test}}\}$. A reliable solver is expected first to identify the representation function and transform the input data $\bm{x}$ to its representation $\Phi^{*}(\bm{x})$. Then it reduces to a regression problem, and some optimization algorithms are performed to find the weight matrix from in-context samples. Finally, it outputs the desired result $\bm{y}_{\text{test}}$ by applying transformations on the test data. We prove by construction that there exists a AlgoFormer that solves the task, akin to the human-designed reliable solver.
@@
[…]
Remarks. The detailed proof is available in Appendix A.1. Our construction of the transformer block involves three distinct sub-transformers, each assigned specific responsibilities. The pre-transformer, characterized by identity attention, is dedicated to representation transformation through feed-forward neural networks. This stage reduces the task to a multivariate regression problem. Subsequently, the looped transformer operates in-context to determine the optimal weight, effectively acting as an iterative solver. Finally, the post-transformer is responsible for the post-processing and generate the desired result $\bm{A}\Phi^{*}\left(\bm{x}_{\text{test}}\right)$. Here, the input prompt to the transformer is formulated as
$\bm{P}=\left[\begin{array}[]{ccccccc}\bm{x}_{1}&\bm{0}&\cdots&\bm{x}_{N}&\bm{0}&\bm{x}_{\text{test}}\\ \bm{0}&\bm{y}_{1}&\cdots&\bm{0}&\bm{y}_{N}&\bm{0}\\ \bm{p}_{1}^{x}&\bm{p}_{1}^{y}&\cdots&\bm{p}_{N}^{x}&\bm{p}_{N}^{y}&\bm{p}_{N+1}^{x}\end{array}\right],$
where $\bm{p}_{i}^{x}$, and $\bm{p}_{i}^{y}$ denote positional embeddings and will be specified in the proof. Due to the differing dimensions of the input $\bm{x}$ and its corresponding label $\bm{y}$, zero padding is incorporated to reshape them into vectors of the same dimension. The structure of the prompt $\bm{P}$ aligns with similar formulations in previous works (Bai et al., 2023; Akyürek et al., 2023; Garg et al., 2022). For different input prompts $\bm{P}$, the hidden linear weights $\bm{A}$ are distinct but the representation function $\Phi^{*}(\cdot)$ is fixed. In comparison with the standard transformer adopted in Guo et al. (2024), which investigates similar tasks, the designed AlgoFormer has a significantly lower parameter size, making it closer to the envisioned human-designed algorithm. Notably, we construct the looped transformer to perform gradient descent for the multi-variate regression. However, the transformer exhibits remarkable versatility, as it has the capability to apply (ridge) regularized regression and more effective optimization algorithms beyond gradient descent. For more details, please refer to Section 4.

### v2

# AlgoFormer: An Efficient Transformer Framework with Algorithmic Structures

@@ Abstract
Besides natural language processing, transformers exhibit extraordinary performance in solving broader applications, including scientific computing and computer vision. Previous works try to explain this from the expressive power and capability perspectives that standard transformers are capable of performing some algorithms. To empower transformers with algorithmic capabilities and motivated by the recently proposed looped transformer (Yang et al., 2024; Giannou et al., 2023), we design a novel transformer framework, dubbed Algorithm Transformer (abbreviated as AlgoFormer). We provide an insight that efficient transformer architectures can be designed by leveraging prior knowledge of tasks and the underlying structure of potential algorithms. Compared with the standard transformer and vanilla looped transformer, the proposed AlgoFormer can perform efficiently in algorithm representation in some specific tasks. In particular, inspired by the structure of human-designed learning algorithms, our transformer framework consists of a pre-transformer that is responsible for task preprocessing, a looped transformer for iterative optimization algorithms, and a post-transformer for producing the desired results after post-processing. We provide theoretical evidence of the expressive power of the AlgoFormer in solving some challenging problems, mirroring human-designed algorithms. Furthermore, some theoretical and empirical results are presented to show that the designed transformer has the potential to perform algorithm representation and learning. Experimental results demonstrate the empirical superiority of the proposed transformer in that it outperforms the standard transformer and vanilla looped transformer in some specific tasks. An extensive experiment on real language tasks (e.g., neural machine translation of German and English, and text classification) further validates the expressiveness and effectiveness of AlgoFormer.
@@ 1 Introduction
The emergence of the transformer architecture (Vaswani et al., 2017) marks the onset of a new era in natural language processing. Transformer-based large language models (LLMs), such as BERT (Devlin et al., 2019) and GPT-3 (Brown et al., 2020), revolutionized impactful language-centric applications, including language translation (Vaswani et al., 2017; Raffel et al., 2020), text completion/generation (Radford et al., 2019; Brown et al., 2020), sentiment analysis (Devlin et al., 2019), and mathematical reasoning (Imani et al., 2023; Yu et al., 2024). Beyond the initial surge in LLMs, these transformer-based models have found extensive applications in diverse domains such as computer vision (Dosovitskiy et al., 2021), time series (Li et al., 2019), bioinformatics (Zhang et al., 2023b), and addressing various physical problems (Cao, 2021). While many studies have concentrated on employing transformer-based models to tackle challenging real-world tasks, yielding superior performances compared to earlier models, the mathematical understanding of transformers remains incomplete.
@@
Garg et al. (2022) empirically investigate the performance of transformers in in-context learning, where the input tokens are input-label pairs generated from classical machine learning models, e.g., (sparse) linear regression and decision tree. They find that transformers can perform comparably as standard human-designed machine learning algorithms. Some subsequent works try to explain the phenomenon. Akyürek et al. (2023) characterize decoder-based transformer as employing stochastic gradient descent for linear regression. Bai et al. (2023) demonstrate that transformers can address statistical learning problems and employ algorithm selection, such as ridge regression, Lasso, and classification problems. Zhang et al. (2023a) and Huang et al. (2023) simplify the transformer model with reduced active parameters, yet reveal that the simplified transformer retains sufficient expressiveness for in-context linear regression problems. In Ahn et al. (2023), transformers are extended to implement preconditioned gradient descent. The looped transformer is proposed in Giannou et al. (2023), and is shown to have the potential to perform basic operations (e.g., addition and multiplication), as well as implicitly learn iterative algorithms  Yang et al. (2024). More related and interesting studies can be found in Huang et al. (2023); Von Oswald et al. (2023); Mahankali et al. (2024).
@@
In this paper, inspired by the recently proposed looped transformer (Yang et al., 2024; Giannou et al., 2023), we propose a novel transformer framework, which we refer to as AlgoFormer, and strictly enforce it as an algorithm learner by the structure regularization on its architecture. The transformer framework consists of three modules (sub-transformers), i.e., the pre-, looped, and post-transformers, designed to perform distinct roles. The pre-transformer is responsible for preprocessing the input data, and formulating it into some mathematical problems. The looped transformer acts as an iterative algorithm in solving the hidden problems. Finally, the post-transformer handles suitable postprocessing to produce the desired results. In contrast to standard transformers, the AlgoFormer is more likely to implement algorithms, due to its algorithmic structures shown in Figure 1.
2 Motivation and Proposed Method
@@
In this section, we mainly discuss the construction and intuition of the AlgoFormer. Its advantages over the standard transformer is then conveyed. Before going into details, we first elaborate the mathematical definition of transformer layers.
@@ 2.1 Preliminaries
A one-layer transformer is mathematically formulated as:
@@ 2.2 Algorithmic Structures of Transformers
Figure 1: Algorithmic structure of the AlgoFormer. Here, $\text{TF}_{\text{pre}}$, $\text{TF}_{\text{loop}}$, and $\text{TF}_{\text{post}}$ are multi-layer transformers; “statements” represent some fundamental operations in classical algorithms.
As discussed in the introduction, rather than simply interpreting it as an implicit function approximator, the transformer may in-context execute some implicit algorithms learned from training data. However, it is still unverified whether the standard multi-layer transformer is exactly performing algorithms.
AlgoFormer. As shown in the green part of Figure 1, vanilla looped transformers (Yang et al., 2024; Giannou et al., 2023) admit the same structure as iterative algorithms. However, real applications are usually much more complicated. For example, given a task with data pairs, a well-trained researcher may first pre-process the data under some prior knowledge, and then formulate a mathematical (optimization) problem. Following that, some designed solvers, usually iterative algorithms, are performed. Finally, the desired results are obtained after further post-processing. The designed AlgoFormer (Algorithm Transformer), visualized in Figure 1, enjoys the same structure as wide classes of algorithms. Specifically, we separate the transformer framework into three parts, i.e., pre-transformer $\text{TF}_{\text{pre}}$, looped transformer $\text{TF}_{\text{loop}}$, and post-transformer $\text{TF}_{\text{post}}$. Here, those three sub-transformers are standard multi-layer transformers in Equation (1). Given the input token vectors $\bm{X}$ and the number of iteration steps $T$, the output admits:
@@
The looped transformer layers ($\text{TF}_{\text{loop}}$) share the same set of weights, as they perform identical computations during each iteration. In contrast, the pre-transformer ($\text{TF}_{\text{pre}}$) and post-transformer ($\text{TF}_{\text{post}}$) utilize distinct weights and architectures to handle their specific computational roles. The hyperparameters for each transformer module (e.g., number of heads, layers, and hidden dimensions) are configured based on prior knowledge of the computational complexity of the target algorithms. Although we present AlgoFormer as in Equation (2), which consists of three modules, the key insight here is that the AlgoFormer can be designed flexibly based on the prior knowledge of the algorithm structure for the given task. Importantly, we do not restrict AlgoFormer to the specific form outlined in Equation (2). Compared with standard transformers, the AlgoFormer acts more as the algorithm learner, by strictly regularizing the loop structure. In contrast to Giannou et al. (2023) and Yang et al. (2024), which primarily focus on tasks solvable by iterative algorithms, our approach introduces additional transformer modules (e.g., pre- and post-transformers) for processing. These components are crucial for addressing the processing needs of real-world applications. This design makes the AlgoFormer capable of representing more complex algorithms and solving challenging tasks more efficiently. Additionally, one of the core insights of our work is that transformer architectures can be designed more efficiently, flexibly, and diversely by leveraging prior knowledge and the pre-defined structure of potential algorithms. This approach enables AlgoFormer to generalize across a broader range of applications while maintaining high efficiency and adaptability, such as representing algorithms involving nested loops, multiple loops, or multi-processing.
2.3 Training Strategy
Our training strategy builds upon the methodology introduced in Yang et al. (2024). Let $\bm{P}^{i}=[\bm{x}_{1},f(\bm{x}_{1}),\cdots,\bm{x}_{i-1},f(\bm{x}_{i-1}),\bm{x}_{i}]$ represents the input prompt for $1\leq i\leq N$. We denote the AlgoFormer as $\text{TF}_{\text{Algo}}^{t}(\cdot;\bm{\Theta})$, where $f(\cdot)$ is a task-specific function that varies across different sequences and $t$ indicates the number of loops (iterations) in Equation 2, and $\bm{\Theta}$ represent the transformer parameters. Instead of evaluating the loss solely on $\text{TF}_{\text{Algo}}^{T}(\cdot;\bm{\Theta})$ with $T$ iterations, we minimize the expected loss over averaged iteration numbers:
where $T_{0}=\max\{T-\Delta T,0\}$ is the initial step for evaluating performance, $T$ is the maximal loop iterations during training, and $\Delta T$ is the number of loop iterations included in the loss. To stabilize the training, we adopt the moving average over loop iterations from $T_{0}$ to $T$ in the loss. Incorporating the loss from $T_{0}$ to $T$ ensures that the transformer module faithfully applies iterative algorithms and generalizes beyond the number of training loops. This approach helps enhance the transformer’s capability for tasks requiring a varying number of iterations. Experimental results in Sections 5.2 and 5.3 further demonstrate that the transformer not only applies certain iterative algorithms but also generalizes well to longer loop iterations beyond those encountered during training.
@@ 3 Expressive Power
In this section, we theoretically show by construction that AlgoFormer is capable of solving some challenging tasks, akin to human-designed algorithms. The core idea is as follows. Initially, the pre-transformer undertakes the crucial task of preprocessing the input data, such as representation transformation. The looped transformer is responsible for iterative algorithms in optimization problems. Finally, it is ready to output the desired result by the post-transformer. Through the analysis of AlgoFormer’s expressive power in addressing these tasks, we expect its potential to make contributions to the communities of scientific computing and machine learning. Throughout the section, we assume that the maximal number of data samples is $N$ and all sample observations (e.g., $\bm{x}_{i}$ and $\bm{y}_{i}$) are bounded.
@@ 3.1 Regression with Representation
We consider regression problems with representation, where the output behaves as a linear function of the input with a fixed representation function. Here, we adopt the $L$-layer MLPs with (leaky) ReLU activation function as the representation function $\Phi^{*}(\cdot)$. Specifically, we generate each in-context sample by first sampling the linear weight $\bm{A}$ from the prior $\mathcal{P}_{A}$, and then generating the input-label pair $\{(\bm{x}_{i},\bm{y}_{i})\}$ with $\bm{x}_{i}\in\mathbb{R}^{d}\sim\mathcal{P}_{x}$, $\bm{y}_{i}=\bm{A}\Phi^{*}(\bm{x}_{i})+\bm{\epsilon}_{i}$ and $\bm{\epsilon}_{i}\sim\mathcal{N}\left(\bm{0},\sigma^{2}\bm{I}\right)$. We aim to find the test label $\bm{y}_{\text{test}}:=\bm{A}\Phi^{*}(\bm{x}_{\text{test}})$, given the in-context samples and test data $\{\bm{x}_{1},\bm{y}_{1},\cdots,\bm{x}_{N},\bm{y}_{N},\bm{x}_{\text{test}}\}$. Here, the weight matrix $A$ varies across different sequences of in-context samples but remains constant within a single sequence, and is learned in-context. The representation function $\Phi^{*}$, on the other hand, is fixed across all samples of sequences and is learned during training. A reliable solver is expected first to identify the representation function and transform the input data $\bm{x}$ to its representation $\Phi^{*}(\bm{x})$. Then it reduces to a regression problem, and some optimization algorithms are performed to find the weight matrix from in-context samples. Finally, it outputs the desired result $\bm{y}_{\text{test}}$ by applying transformations on the test data. We prove by construction that there exists an AlgoFormer that solves the task, akin to the human-designed reliable solver.
@@
There exists a designed AlgoFormer with $\text{TF}_{\text{pre}}$ (an $(L+1)$-layer two-head transformer), $\text{TF}_{\text{loop}}$ (a one-layer two-head transformer), and $\text{TF}_{\text{post}}$ (a one-layer one-head transformer), that outputs $\bm{A}\Phi^{*}\left(\bm{x}_{\text{test}}\right)$ from the input-label pairs $\{\bm{x}_{1},\bm{y}_{1},\cdots,\bm{x}_{N},\bm{y}_{N},\bm{x}_{\text{test}}\}$ by fitting the representation function and applying gradient descent for multi-variate regression. The emulation of each step is not exact, as there is some error introduced in each step. However, the error can be made arbitrarily close to zero by increasing the temperature of the softmax and adjusting another free parameter, neither of which affects the size of the network.
Remarks. The detailed proof is available in Appendix A.1. Our construction of the transformer framework involves three distinct sub-transformers, each assigned specific responsibilities. The pre-transformer, characterized by identity attention, is dedicated to representation transformation through feed-forward neural networks. This stage reduces the task to a multivariate regression problem. Subsequently, the looped transformer operates in-context to determine the optimal weight, effectively acting as an iterative solver. Finally, the post-transformer is responsible for the post-processing and generate the desired result $\bm{A}\Phi^{*}\left(\bm{x}_{\text{test}}\right)$. Here, the input prompt to the transformer is formulated as $\bm{P}=\left[\begin{array}[]{ccccccc}\bm{x}_{1}&\bm{0}&\cdots&\bm{x}_{N}&\bm{0}&\bm{x}_{\text{test}}\\ \bm{0}&\bm{y}_{1}&\cdots&\bm{0}&\bm{y}_{N}&\bm{0}\\ \bm{p}_{1}^{x}&\bm{p}_{1}^{y}&\cdots&\bm{p}_{N}^{x}&\bm{p}_{N}^{y}&\bm{p}_{N+1}^{x}\end{array}\right],$ where $\bm{p}_{i}^{x}$, and $\bm{p}_{i}^{y}$ denote positional embeddings and will be specified in the proof. Due to the differing dimensions of the input $\bm{x}$ and its corresponding label $\bm{y}$, zero padding is incorporated to reshape them into vectors of the same dimension. The structure of the prompt $\bm{P}$ aligns with similar formulations in previous works (Bai et al., 2023; Akyürek et al., 2023; Garg et al., 2022). For different input prompts $\bm{P}$, the hidden linear weights $\bm{A}$ are distinct but the representation function $\Phi^{*}(\cdot)$ is fixed. In comparison with the standard transformer adopted in Guo et al. (2024), which investigates similar tasks, the designed AlgoFormer has a significantly lower parameter size, making it closer to the envisioned human-designed algorithm. Notably, we construct the looped transformer to perform gradient descent for the multi-variate regression. However, the transformer exhibits remarkable versatility, as it has the capability to apply (ridge) regularized regression and more effective optimization algorithms beyond gradient descent. For more details, please refer to Section 4.